40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4145     Entry time: Wed Jan 12 22:19:54 2011
Author: Koji 
Type: Update 
Category: IOO 
Subject: MC flakiness solved 

[Koji Suresh Kiwamu]

Suresh modified the MC board to have +5V offset at the output. (To be reported in the separated elog)

The MC lock has not been obtained at this point. An investigation revealed that there was very small (~5mVpp) PDH signal.

Kiwamu removed his triple resonant adapter and put the 50Ohm termination insted.
This restored the signal level normal although this changed the demodulation phase about 20deg.
We left the demodulation phase as it is because this is a temporary setup and the loss of the signal is not significant.

Now the MC is steadily locked with the single super boost.

ELOG V3.1.3-